### v1

# Offline Robotic World Model: Learning Robotic Policies without a Physics Simulator

Abstract
Reinforcement Learning (RL) has demonstrated impressive capabilities in robotic control but remains challenging due to high sample complexity, safety concerns, and the sim-to-real gap. While offline RL eliminates the need for risky real-world exploration by learning from pre-collected data, it suffers from distribution shift, limiting policy generalization. Model-Based RL (MBRL) addresses this by leveraging predictive models for synthetic rollouts, yet existing approaches often lack robust uncertainty estimation, leading to compounding errors in offline settings. We introduce Offline Robotic World Model (RWM-O), a model-based approach that explicitly estimates epistemic uncertainty to improve policy learning without reliance on a physics simulator. By integrating these uncertainty estimates into policy optimization, our approach penalizes unreliable transitions, reducing overfitting to model errors and enhancing stability. Experimental results show that RWM-O improves generalization and safety, enabling policy learning purely from offline data and advancing scalable, data-efficient RL for robotics.
Keywords: Offline Reinforcement Learning, Model-Based Reinforcement Learning, Uncertainty Estimation
Figure 1: Overview of RWM-O and MOPO-PPO. RWM-O is trained on offline data to predict long-horizon dynamics with an ensemble-based uncertainty estimate. The shade visualizes epistemic uncertainty, which is penalized during policy optimization to avoid overfitting to unreliable predictions. Actions are omitted for clarity. Our framework enables direct policy learning without simulators, bridging the gap between model-based offline RL and real-world robotic deployment.
1 Introduction
Reinforcement Learning (RL) has demonstrated significant success in robotic control, enabling policies that outperform traditional control techniques in complex tasks [1, 2, 3, 4]. However, applying RL directly to real-world robotics remains challenging due to the high sample complexity and the risks associated with exploratory interactions [5, 6]. A common alternative is training policies in high-fidelity physics simulators, which offer a safe and efficient environment for data collection and optimization [7, 8, 9]. However, simulators struggle to capture real-world dynamics with perfect accuracy, leading to the well-known sim-to-real gap, which degrades policy performance when transferred to physical hardware [10, 11]. To address these challenges, offline RL has emerged as a promising paradigm by enabling policy learning entirely from pre-collected real-world data, eliminating the need for risky online exploration [12, 13, 14, 15]. Despite its promise, offline RL is prone to distribution shift, where the learned policy encounters states that are insufficiently represented in the dataset, leading to extrapolation errors and poor generalization [13, 14, 12]. Model-free offline RL methods mitigate this by enforcing strict constraints to keep the policy within the dataset’s support, but at the cost of limiting generalization beyond observed transitions [16, 17, 18].
Model-Based Reinforcement Learning (MBRL) provides an attractive alternative by learning a predictive model of environment dynamics, allowing the generation of synthetic experience for policy optimization [19, 20]. This is particularly advantageous in offline RL for three reasons [15]. First, model-based methods benefit from denser supervision, as the dynamics model is trained on all dataset transitions, even when rewards are sparse [21]. Second, training the model via supervised learning provides more stable gradients compared to bootstrapped value estimation. Finally, well-established uncertainty estimation techniques from supervised learning can be leveraged to mitigate distribution shift [22, 23, 24, 25]. Recent work on the Robotic World Model (RWM) has demonstrated the potential of neural network-based dynamics models for policy learning [26]. RWM effectively captures long-horizon predictive dynamics without domain-specific inductive biases, enabling robust policy optimization through Model-Based Policy Optimization with Proximal Policy Optimization (MBPO-PPO) [27]. However, existing formulations lack explicit uncertainty estimation, making them susceptible to distribution shift in offline settings. Without uncertainty awareness, RWM requires online corrections to mitigate errors introduced by out-of-distribution states [12, 28, 20], which limits its applicability in fully offline RL [29, 30].
In this work, we extend RWM by incorporating uncertainty-aware dynamics modeling to enable robust policy learning in offline RL. We propose an ensemble-based architecture that explicitly estimates model uncertainty, allowing the policy to incorporate uncertainty-aware rollouts. To leverage this information, we adapt Model-Based Offline Policy Optimization (MOPO) [15] to PPO, enabling safer policy learning by penalizing high-uncertainty state transitions without requiring additional real-world interactions. Our approach enables policy learning purely from offline robotic data, eliminating the need for a physics simulator or online interactions. To our knowledge, our locomotion experiments mark the first demonstration of a locomotion policy learned entirely from offline data in a model-based fashion, with successful deployment on hardware. By enhancing the reliability and adaptability of MBRL, our approach contributes toward scalable, data-efficient policy learning, facilitating broader adoption of offline RL in real-world robotics. Supplementary videos for this work are available at https://sites.google.com/view/corl2025-rwm-o/home.
2 Related work
2.1 Model-Based Reinforcement Learning
Model-Based Reinforcement Learning (MBRL) constructs an explicit predictive model of environment dynamics, improving sample efficiency and facilitating long-horizon planning. Unlike model-free RL, which directly learns policies from data, MBRL generates synthetic rollouts to optimize policies, reducing reliance on real-world interactions. Early approaches such as PILCO leverage Gaussian processes to model system dynamics, achieving data-efficient learning but facing scalability issues in high-dimensional spaces [12]. Local linear models, as seen in guided policy search [28], improve tractability in complex control tasks [31]. Neural network-based dynamics models have since become the dominant approach, allowing for greater expressivity and scalability [21, 32]. One notable example is Model-Based Policy Optimization (MBPO), which interleaves model-generated rollouts with model-free updates, selectively trusting the learned model to improve stability [27].
Recent advancements in latent-space world models have further improved MBRL. PlaNet introduces a compact latent dynamics model for planning directly in an abstract space, making it computationally feasible for long-horizon tasks [19]. Dreamer extends this idea by integrating an actor-critic framework, achieving strong performance in continuous control [20, 33, 34]. A major challenge in MBRL is model fidelity and its impact on policy learning. Robotic World Model (RWM) addresses this by introducing a self-supervised, autoregressive neural simulator that learns long-horizon environment dynamics without domain-specific inductive biases [26]. Unlike previous methods, RWM enables stable policy optimization via MBPO-PPO, demonstrating successful deployment in real-world robotic systems.
2.2 Uncertainty Estimation and Distribution Shift Mitigation
Incorporating uncertainty estimation into MBRL is crucial, especially in offline settings where the lack of online interactions limits the ability to correct model inaccuracies. One of the most effective and widely used techniques for model uncertainty estimation is the use of bootstrap ensembles, which have been successfully applied in various MBRL frameworks to mitigate overestimation caused by distribution shifts in offline RL [23, 13, 15]. Probabilistic Ensembles with Trajectory Sampling (PETS) pioneers the use of probabilistic ensembles in MBRL, improving robustness by leveraging multiple network predictions for decision-making [13]. To further improve uncertainty estimation and its integration into policy learning, several advanced techniques have been developed. MOUP introduces ensemble dropout networks for uncertainty estimation while incorporating a maximum mean discrepancy constraint into policy optimization, ensuring bounded state mismatch and improving policy learning in offline settings [35]. SUMO characterizes uncertainty by measuring the cross-entropy between model dynamics and true dynamics using a k-nearest neighbor search method [36]. This technique provides an alternative perspective on estimating uncertainty beyond conventional model ensembles.
Uncertainty-aware policy optimization is key to mitigating the risks of distribution shift. MORPO explicitly quantifies the uncertainty of predicted dynamics and incorporates a reward penalty to balance optimism and conservatism in policy learning [37]. Similarly, COMBO tackles uncertainty estimation challenges by enforcing conservatism in value estimation without relying on explicit uncertainty quantification [38]. Our work extends these ideas by integrating uncertainty estimation into RWM and adapting MOPO to PPO, enhancing stability and robustness in offline robotic control.
3 Preliminaries
3.1 Offline Model-Based Reinforcement Learning
We formulate the problem by modeling the environment as a Partially Observable Markov Decision Process (POMDP) [39], defined by the tuple $\left({\mathcal{S}},{\mathcal{A}},{\mathcal{O}},T,R,O,\gamma\right)$, where ${\mathcal{S}}$, ${\mathcal{A}}$, and ${\mathcal{O}}$ denote the state, action, and observation spaces, respectively. The transition kernel $T:{\mathcal{S}}\times{\mathcal{A}}\to{\mathcal{S}}$ captures the environment dynamics, while the reward function $R:{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}\to{\mathbb{R}}$ maps transitions to scalar rewards. The agent seeks to learn a policy $\pi_{\theta}:{\mathcal{O}}\to{\mathcal{A}}$ that maximizes the expected discounted return $\mathbb{E}_{\pi_{\theta}}\left[\sum_{t\geq 0}\gamma^{t}r_{t}\right]$, where $r_{t}$ is the reward at time $t$ and $\gamma\in\left[0,1\right]$ is the discount factor.
In offline RL, the agent is trained using a fixed dataset of transitions collected by one or a mixture of behavior policies. Unlike online RL, the agent cannot interact with the environment, making it susceptible to extrapolation errors when encountering out-of-distribution states [29, 30]. Many offline RL methods impose conservatism by constraining the policy to remain within the dataset’s support, limiting generalization beyond observed trajectories [17, 18]. MBRL provides an alternative approach by learning a world model of the environment, which can be leveraged for policy optimization in imagination [40]. Instead of relying solely on offline data for policy learning, MBRL methods construct a predictive model of the transition dynamics, allowing the agent to generate synthetic rollouts and expand the effective training data. This enables more efficient learning and better generalization, which is particularly important in offline settings [15]. However, model errors can accumulate over long-horizon predictions, introducing bias and degrading policy performance. Addressing these challenges requires accurate long-horizon modeling and robust handling of model errors.
3.2 Robotic World Model
Robotic World Model (RWM) is a neural network-based world model designed to accurately capture long-horizon robotic dynamics without requiring domain-specific inductive biases [26]. Unlike conventional methods that rely on structured priors or handcrafted features, RWM generalizes across diverse robotic systems using a self-supervised, dual-autoregressive training framework. RWM learns to predict future observations autoregressively, feeding its own predictions back into the model to simulate long-horizon rollouts. The training objective minimizes a multi-step prediction loss
${\mathcal{L}}=\frac{1}{N}\sum_{k=1}^{N}\alpha^{k}\left[L_{o}\left(o^{\prime}_{t+k},o_{t+k}\right)+L_{c}\left(c^{\prime}_{t+k},c_{t+k}\right)\right],$ (1)
where $N$ denotes the autoregressive prediction steps, $o^{\prime}_{t+k}$ and $c^{\prime}_{t+k}$ are the predicted observations and privileged information (e.g., contact information), $L_{o}$ and $L_{c}$ measure the prediction discrepancy, and $\alpha_{k}$ is a decay factor. The predicted observation $k$ steps ahead of time $t$ can be written as
$o^{\prime}_{t+k}\sim p_{\phi}\left(\cdot\mid o_{t-M+k:t},o^{\prime}_{t+1:t+k-1},a_{t-M+k:t+k-1}\right),$ (2)
where $M$ denotes the history horizon steps. Through this design, RWM achieves robust trajectory forecasting across manipulation and locomotion tasks. Policies trained with RWM have been zero-shot deployed on hardware, achieving high-fidelity control without requiring additional fine-tuning.
Model-Based Policy Optimization (MBPO) trains policies from learned world models by interleaving real environment data with simulated imagination rollouts [27]. RWM extends MBPO to PPO with extensive on-policy updates, resulting in MBPO-PPO, a model-based RL framework tailored for long-horizon robotic control [26]. MBPO-PPO follows the standard MBPO approach but benefits from high-fidelity RWM rollouts, which significantly improve stability and accuracy during policy learning.
4 Approach
4.1 Uncertainty Quantification with Offline Robotic World Model
In MBRL, the learned dynamics model inevitably becomes inaccurate when making predictions for state-action pairs that deviate from the distribution of the training data. This issue is particularly critical in the offline setting, where errors in the learned dynamics cannot be corrected through additional environment interactions. As a result, standard model-based policy optimization methods risk overfitting to erroneous model predictions in out-of-distribution regions, leading to suboptimal or unsafe policy behavior. To ensure reliable performance, it is essential to balance the trade-off between policy improvement and model reliability: leveraging model-based imagination to explore high-return policies beyond the support of the dataset while mitigating the risk of policy degradation due to overfitting to model errors in regions with high uncertainty.
To this end, we introduce Offline Robotic World Model (RWM-O), where we explicitly incorporate uncertainty quantification into the dynamics model based on RWM. We aim to design an uncertainty estimator that captures both epistemic uncertainty, which arises due to limited training data, and aleatoric uncertainty, which reflects inherent stochasticity in the environment. Bootstrap ensembles have been shown to provide consistent estimates of uncertainty [41] and have been successfully applied in MBRL to improve robustness [13]. RWM-O extends RWM by introducing an ensemble-based uncertainty-aware architecture. Specifically, we apply bootstrap ensembles to the observation prediction head of the model after a shared recurrent feature extractor (e.g., a GRU). Each ensemble element $b$ independently predicts the mean and variance of a Gaussian distribution over the next observation
$o^{\prime b}_{t+k}\sim{\mathcal{N}}\left(\mu_{o_{t+k}}^{b},\sigma_{o_{t+k}}^{2,b}\right),$ (3)
where $\mu_{o_{t+k}}^{b},\sigma_{o_{t+k}}^{2,b}$ denote the mean and variance predicted by ensemble member $b$, following the autoregressive prediction scheme in Eq. 2. The learned variance component captures aleatoric uncertainty effectively. The training objective is then computed as the average loss over all ensemble members, following Eq. 1.
During inference, the ensemble mean is used as the predicted next observation, while the variance across ensemble members provides an estimate of epistemic uncertainty $u_{p_{\phi}}$, depending on the world model $p_{\phi}$. The training overview is visualized in Fig. 1.
$o^{\prime}_{t+1}=\mathbb{E}_{b}\left[\mu_{o_{t+1}}^{b}\right],u_{t+1}=u_{p_{\phi}}\left(o_{t-M+1:t},a_{t-M+k:t}\right)=\mathrm{Var}_{b}\left[\mu_{o_{t+1}}^{b}\right].$ (4)
By incorporating uncertainty-aware modeling into the dynamics learning process, RWM-O enables robust trajectory forecasting in offline settings. The explicit quantification of epistemic uncertainty allows for uncertainty-informed policy optimization, preventing the policy from over-relying on uncertain model predictions.
4.2 Policy Optimization on Uncertainty-Penalized Imagination
To ensure robust policy learning in offline MBRL, we introduce Model-Based Offline Policy Optimization with Proximal Policy Optimization (MOPO-PPO), an adaptation of the MOPO framework [15] to PPO [42]. MOPO-PPO allows cautious exploration beyond the behavioral distribution while penalizing transitions with high epistemic uncertainty. Specifically, given the uncertainty estimator $u_{p_{\phi}}$ from Eq. 4, we modify the reward function to discourage the policy from exploiting unreliable model predictions:
$\tilde{r}\left(o_{t},a_{t}\right)=r_{t}\left(o_{t},a_{t}\right)-\lambda u_{p_{\phi}}\left(o_{t-M+1:t},a_{t-M+k:t}\right),$ (5)
where $\lambda$ is a hyperparameter controlling the penalty strength. This formulation encourages the policy to focus on high-confidence regions of the learned dynamics model while avoiding overfitting to model errors.
The training process of MOPO-PPO using RWM-O is outlined in Fig. 1 and Algorithm 1.
Algorithm 1 Offline policy optimization with RWM-O
1:  Input: Offline dataset ${\mathcal{D}}$
2:  Initialize policy $\pi_{\theta}$, world model $p_{\phi}$
3:  Train the world model $p_{\phi}$ autoregressively using offline dataset ${\mathcal{D}}$ according to Eq. 1
4:  for learning iterations $=1,2,\dots$ do
5:     Initialize imagination agents with observations sampled from ${\mathcal{D}}$
6:     Roll out uncertainty-aware trajectories using $\pi_{\theta}$ and $p_{\phi}$ for $T$ steps according to Eq. 4
7:     Compute uncertainty-penalized rewards for each imagination transition according to Eq. 5
8:     Update $\pi_{\theta}$ using PPO or another RL algorithm
9:  end for
By integrating uncertainty-aware dynamics modeling with policy optimization, RWM-O and MOPO-PPO enable safer and more reliable offline MBRL. Explicit uncertainty estimation mitigates distribution shift, allowing cautious generalization beyond the dataset while ensuring stability through trust-region optimization. By leveraging uncertainty-aware rollouts, MOPO-PPO enhances policy robustness and efficiency, advancing scalable offline RL for real-world robotics.
5 Experiments
We evaluate RWM-O and MOPO-PPO through a comprehensive set of experiments spanning diverse robotic platforms, tasks, and dataset sources. Our objective is to assess the capability of RWM-O in quantifying uncertainty during long-horizon rollouts and to demonstrate the effectiveness of MOPO-PPO in improving policy performance through uncertainty-aware optimization.
5.1 Autoregressive Uncertainty Estimation
A reliable estimate of epistemic uncertainty is critical for ensuring robust policy learning in model-based offline RL. In particular, the ability of a world model to quantify its own uncertainty over long-horizon rollouts determines the trustworthiness of its synthetic experience. To evaluate this property in RWM-O, we analyze the alignment between its predicted epistemic uncertainty and actual model prediction errors during autoregressive trajectory forecasting on ANYmal D. The observation and action spaces of the world model are detailed in Table S1 and Table S3.
We train the world model on offline data collected from a velocity-tracking policy, operating at a control frequency of $50\,Hz$. The model is trained with a history horizon of $M=32$ and a prediction horizon of $N=8$, using an ensemble of five networks to estimate epistemic uncertainty. To evaluate uncertainty estimation, we compute epistemic uncertainty as the variance of the ensemble predictions, following Eq. 4. Aleatoric uncertainty is estimated as the mean of the predicted standard deviations across ensemble members. Details of the network architecture and training setup are provided in Sec. A.2.1 and Sec. A.3.1. The predictive error, epistemic uncertainty, and aleatoric uncertainty over autoregressive rollouts are visualized in Fig. 2.
Figure 2: Autoregressive trajectory prediction (left) and uncertainty estimation (right) by RWM-O. Solid lines represent ground truth trajectories, while dashed lines denote predicted state evolution. Predictions commence at $t=32$ using historical observations, with future observations predicted autoregressively by feeding prior predictions back into the model.
At $t=32$, the model transitions from conditioning on ground-truth observations to fully autoregressive rollouts, wherein predicted states are recursively fed back into the model. As expected, the accumulated prediction error (grey) increases over time due to compounding model inaccuracies. Importantly, the estimated epistemic uncertainty (dark blue) closely follows the trend of the prediction error, demonstrating that RWM-O effectively captures uncertainty in regions where the model generalization deteriorates. In contrast, the aleatoric uncertainty (light blue) remains low, reflecting small stochasticity in the environment. The strong correlation between epistemic uncertainty and model prediction error justifies its role as a trust metric for policy optimization. By penalizing high-uncertainty transitions, MOPO-PPO prevents policy exploitation in poorly modeled regions of the state space, mitigating the risk of compounding errors in long-horizon rollouts. These results validate the proposed uncertainty-aware regularization strategy, reinforcing the efficacy of RWM-O in enabling robust policy learning in offline RL.
5.2 Uncertainty-Penalized Imagination
To evaluate the role of uncertainty-aware regularization in policy learning, we train RWM-O on an offline dataset collected from an expert velocity-tracking policy on ANYmal D. We then optimize policies with MOPO-PPO using different values of the uncertainty penalty coefficient $\lambda$ in Eq. 5, which controls the trade-off between exploration and model reliability. Figure 3 shows the evolution of the imagination reward and epistemic uncertainty over training iterations, as well as the final policy evaluation performance.
Figure 3: Imagination reward (left) and epistemic uncertainty (right) during MOPO-PPO training. The policy evaluation on real dynamics is visualized in dots (left). Small penalties lead to overconfident but unreliable policies, while large penalties result in overly conservative behaviors. A well-calibrated penalty (dark blue) achieves the exploration-exploitation balance.
When the penalty is small ($\lambda\leq 0.5$), the policy attains high rewards in imagination but exploits model inaccuracies, leading to overfitting to hallucinated rollouts. This results in poor evaluation performance, as the learned policy exhibits overconfident behaviors in simulation but fails catastrophically when tested on real dynamics. In these cases, the robot often struggles to maintain stable locomotion, exhibiting unintended collisions or failing to walk altogether. In contrast, large penalties ($\lambda\geq 2.0$) force the policy to remain in low-uncertainty regions, limiting exploration and leading to overly conservative behaviors. These policies exhibit excessive caution, often resulting in the robot standing in place or making only minor, ineffective movements. Consequently, the learned policy fails to make meaningful progress, as indicated by the low final evaluation rewards.
With an appropriate penalty ($\lambda=1.0$), the policy effectively balances exploration and exploitation. The imagination reward increases steadily, while epistemic uncertainty remains controlled, allowing the policy to leverage synthetic rollouts without overfitting to model errors. This results in effective learning, where the robot is able to explore new behaviors in imagination while remaining in trustworthy regions of the learned model. The strong alignment between training progress and final evaluation performance highlights the effectiveness of MOPO-PPO in leveraging epistemic uncertainty to mitigate distribution shift, ultimately enhancing policy robustness in offline MBRL.
5.3 Generality across Robotic Environments
To evaluate the generality and robustness of MOPO-PPO with RWM-O across diverse robotic environments, we compare its performance against widely adopted offline RL baselines, including Conservative Q-Learning (CQL) [43] with Soft Actor-Critic (SAC) [44], MBPO [27], MBPO-PPO [26], and MOPO [15]. These baselines represent state-of-the-art approaches in model-free and model-based offline RL. For a fair comparison, we train MBPO and MBPO-PPO in an offline setting, treating them as uncertainty-unaware counterparts to MOPO and MOPO-PPO.
We benchmark these methods on three robotic tasks in simulation: Reach-Franka (manipulation), Velocity-G1 (humanoid locomotion), and Velocity-ANYmal-D (quadruped locomotion), spanning a broad range of dynamics complexities and control challenges. Each method is trained on four dataset types: Random: trajectories collected from a randomly initialized policy, Medium: trajectories from a partially trained PPO policy, Expert: trajectories from a fully trained PPO policy, Mixed: a combination of replay buffer data from PPO training at various stages. All methods are trained and evaluated under identical conditions. The normalized episodic rewards for each method across different environments and dataset types are presented in Fig. 4.
Figure 4: Normalized episodic rewards across diverse robotic environments and offline RL algorithms with different dataset types. MOPO-PPO consistently outperforms uncertainty-unaware baselines, particularly in complex locomotion tasks.
Across all methods, performance improves with increasing dataset quality, as richer data distributions reduce the impact of distribution shift. Policies trained on the Mixed dataset achieve the highest rewards due to its broad state-action coverage, comparable to those trained on Expert data, which tends to overfit to specific dynamics and biases the policy towards suboptimal generalization.
While CQL and SAC-based model-based methods perform competitively in the manipulation task, they demonstrate significantly weaker performance in locomotion tasks, where the complexity of dynamic interactions amplifies the limitations of value-based learning. As highlighted in Sec. 5.2, MBPO and MBPO-PPO, which lack explicit uncertainty estimation, tend to overfit to model inaccuracies, leading to overconfident policies. In manipulation tasks, this overfitting is mitigated when dataset coverage is sufficient. However, in locomotion tasks, it results in catastrophic failures, including robot collisions and falls.
In contrast, MOPO-PPO consistently outperforms baselines across all tasks, particularly in Velocity-G1 and Velocity-ANYmal-D, where uncertainty-aware rollouts significantly enhance policy stability. By penalizing high-uncertainty transitions, MOPO-PPO prevents policies from exploiting hallucinated model predictions, leading to safer and more reliable locomotion behaviors. The deployment of the learned policy is visualized in Fig. S5. The strong performance of MOPO-PPO across different dataset types and environments demonstrates its effectiveness in mitigating distribution shift, reinforcing its applicability to real-world robotic control.
5.4 Offline Learning from Real-World Data
To validate the real-world applicability of RWM-O and MOPO-PPO, we train a locomotion policy entirely from offline data collected on the ANYmal D robot, without any reliance on a physics simulator. We develop an autonomous data collection pipeline with minimal human intervention in Sec. A.4.2. This experiment demonstrates that our approach eliminates the sim-to-real gap by learning directly from real-world experience. By leveraging uncertainty-aware rollouts, MOPO-PPO enables safe and effective policy optimization purely from offline data. A video demonstrating the learned policy running on hardware is available on our project webpage.
6 Conclusion
In this work, we introduce RWM-O, an uncertainty-aware MBRL framework designed to improve offline policy learning. By explicitly estimating epistemic uncertainty through an ensemble-based approach, RWM-O enables uncertainty-aware rollouts, reducing the risk of overfitting to inaccurate model predictions. We further integrate this uncertainty estimation into policy optimization with MOPO-PPO, penalizing unreliable transitions and enhancing policy stability. Our experiments demonstrate that RWM-O improves policy generalization, robustness, and safety across various robotic tasks, outperforming standard model-based and model-free offline RL baselines. Notably, our approach enables policy learning purely from real-world data, eliminating the need for high-fidelity physics simulators and addressing a key limitation of current RL methodologies. By bridging the gap between model-based offline RL and real-world deployment, RWM-O and MOPO-PPO enhance the scalability and data efficiency of RL for robotics, paving the way for more reliable and adaptable policy learning.
Limitations
While RWM-O improves offline policy learning by incorporating uncertainty estimation, several challenges remain. First, the accuracy of the learned dynamics model is inherently limited by the quality and diversity of the offline dataset. If the dataset lacks sufficient coverage of the state-action space, the model may struggle with generalization, leading to unreliable predictions in unobserved regions. Second, although our ensemble-based uncertainty estimation helps mitigate distribution shift, it does not entirely eliminate compounding errors over long-horizon rollouts. The effectiveness of the uncertainty penalty in MOPO-PPO relies on careful tuning of hyperparameters, such as the penalty weight, which may require task-specific adjustments. Finally, while RWM-O and MOPO-PPO remove the need for a physics simulator, the reliance on offline real-world data means the policy cannot explore beyond the dataset distribution. In cases where real-world data is insufficient, supplementing it with simulation data could improve the model’s coverage, but this reintroduces dependence on a physics simulator. This trade-off can limit applicability in scenarios requiring extensive exploration or rapid adaptation to new environments.
Acknowledgments
This research was supported by the ETH AI Center.
Appendix A Appendix
A.1 Task Representation
A.1.1 Observation and action spaces
The observation space for the ANYmal world model is composed of base linear and angular velocities $v$, $\omega$ in the robot frame, measurement of the gravity vector in the robot frame $g$, joint positions $q$, velocities $\dot{q}$ and torques $\tau$ as in Table S1.
Table S1: World model observation space
Entry Symbol Dimensions
base linear velocity $v$ 0:3
base angular velocity $\omega$ 3:6
projected gravity $g$ 6:9
joint positions $q$ 9:21
joint velocities $\dot{q}$ 21:33
joint torques $\tau$ 33:45
The privileged information is used to provide an additional learning objective that implicitly embeds critical information for accurate long-term predictions. The space is composed of knee and foot contacts as in Table S2.
Table S2: World model privileged information space
Entry Symbol Dimensions
knee contact -- 0:4
foot contact -- 4:8
The action space is composed of joint position targets as in Table S3.
Table S3: Action space
Entry Symbol Dimensions
joint position targets $q^{*}$ 0:12
The observation space for the ANYmal velocity tracking policy is composed of base linear and angular velocities $v$, $\omega$ in the robot frame, measurement of the gravity vector in the robot frame $g$, velocity command $c$, joint positions $q$ and velocities $\dot{q}$ as in Table S4.
Table S4: Policy observation space
Entry Symbol Dimensions
base linear velocity $v$ 0:3
base angular velocity $\omega$ 3:6
projected gravity $g$ 6:9
velocity command $c$ 9:12
joint positions $q$ 12:24
joint velocities $\dot{q}$ 24:36
A.1.2 Reward functions
The total reward is sum of the following terms with weights detailed in Table S5.
Table S5: Reward weights
Symbol $w_{v_{xy}}$ $w_{\omega_{z}}$ $w_{v_{z}}$ $w_{\omega_{xy}}$ $w_{q_{\tau}}$
Value $1.0$ $0.5$ $-2.0$ $-0.05$ $-2.5e^{-5}$
Symbol $w_{\ddot{q}}$ $w_{\dot{a}}$ $w_{f_{a}}$ $w_{c}$ $w_{g}$
Value $-2.5e^{-7}$ $-0.01$ $0.5$ $-1.0$ $-5.0$
Linear velocity tracking $x,y$
$r_{v_{xy}}=w_{v_{xy}}e^{-\|c_{xy}-v_{xy}\|_{2}^{2}/\sigma_{v_{xy}}^{2}},$
where $\sigma_{v_{xy}}=0.25$ denotes a temperature factor, $c_{xy}$ and $v_{xy}$ denote the commanded and current base linear velocity.
Angular velocity tracking
$r_{\omega_{z}}=w_{\omega_{z}}e^{-\|c_{z}-\omega_{z}\|_{2}^{2}/\sigma_{\omega_{z}}^{2}},$
where $\sigma_{\omega_{z}}=0.25$ denotes a temperature factor, $c_{z}$ and $\omega_{z}$ denote the commanded and current base angular velocity.
Linear velocity $z$
$r_{v_{z}}=w_{v_{z}}\left\|v_{z}\right\|_{2}^{2},$
where $v_{z}$ denotes the base vertical velocity.
Angular velocity $x,y$
$r_{\omega_{xy}}=w_{\omega_{xy}}\left\|\omega_{xy}\right\|_{2}^{2},$
where $\omega_{xy}$ denotes the current base roll and pitch velocity.
Joint torque
$r_{q_{\tau}}=w_{q_{\tau}}\left\|\tau\right\|_{2}^{2},$
where $\tau$ denotes the joint torques.
Joint acceleration
$r_{\ddot{q}}=w_{\ddot{q}}\left\|\ddot{q}\right\|_{2}^{2},$
where $\ddot{q}$ denotes the joint acceleration.
Action rate
$r_{\dot{a}}=w_{\dot{a}}\|a^{\prime}-a\|_{2}^{2},$
where $a^{\prime}$ and $a$ denote the previous and current actions.
Feet air time
$r_{f_{a}}=w_{f_{a}}t_{f_{a}},$
where $t_{f_{a}}$ denotes the sum of the time for which the feet are in the air.
Undesired contacts
$r_{c}=w_{c}c_{u},$
where $c_{u}$ denotes the counts of the undesired knee contacts.
Flat orientation
$r_{g}=w_{g}g_{xy}^{2},$
where $g_{xy}$ denotes the $xy$-components of the projected gravity.
A.2 Network Architecture
A.2.1 RWM-O
The robotic world model consists of a GRU base and MLP heads predicting the mean and standard deviation of the next observation and privileged information such as contacts, as detailed in Table S6.
Table S6: RWM-O architecture
Component Type Hidden Shape Activation
base GRU 256, 256 --
heads MLP 128 ReLU
A.2.2 MOPO-PPO
The network architectures of the policy and the value function used in MOPO-PPO are detailed in Table S7.
Table S7: Policy and value function architecture
Network Type Hidden Shape Activation
policy MLP 128, 128, 128 ELU
value function MLP 128, 128, 128 ELU
A.3 Training Parameters
The learning networks and algorithm are implemented in PyTorch 2.4.0 with CUDA 12.6 and trained on an NVIDIA RTX 4090 GPU.
A.3.1 RWM-O
The training information of RWM-O is summarized in Table S8.
Table S8: RWM-O training parameters
Parameter Symbol Value
step time seconds $\Delta t$ $0.02$
max iterations -- $2500$
learning rate -- $1e^{-4}$
weight decay -- $1e^{-5}$
batch size -- $1024$
ensemble size $B$ $5$
history horizon $M$ $32$
forecast horizon $N$ $8$
forecast decay $\alpha$ $1.0$
transitions in training data -- $6M$
approximate training hours -- $1$
number of seeds -- $5$
A.3.2 MOPO-PPO
The training information of MOPO-PPO is summarized in Table S9.
Table S9: MOPO-PPO training parameters
Parameter Symbol Value
imagination environments -- $4096$
imagination steps per iteration -- $100$
uncertainty penalty weight $\lambda$ $1.0$
step time seconds $\Delta t$ $0.02$
buffer size $|{\mathcal{D}}|$ $1000$
max iterations -- $2500$
learning rate -- $0.001$
weight decay -- $0.0$
learning epochs -- $5$
mini-batches -- $4$
KL divergence target -- $0.01$
discount factor $\gamma$ $0.99$
clip range $\epsilon$ $0.2$
entropy coefficient -- $0.005$
number of seeds -- $5$
A.4 Offline Policy Learning
A.4.1 Policy Deployment Learned with MOPO-PPO
The policy deployment learning with MOPO-PPO is visualized in Fig. S5.
Figure S5: MOPO-PPO policy deployment learned from offline Mixed dataset across various environments.
A.4.2 Real-World Data Collection
Although RWM-O can be trained on offline data from arbitrary sources, we develop an autonomous pipeline for collecting additional real-world data with minimal human intervention. The robot executes a velocity-tracking policy and records transitions as described in Sec. A.1.1. A visualization of the collection setup is shown in Fig. S6. To ensure safety, a rectangular boundary is defined relative to the odometry origin. Robot odometry is computed by fusing legged odometry and LiDAR data. At each sampling interval, a base velocity command (green arrow) is randomly drawn, and the resulting final pose (red sphere) is predicted via finite integration. If this predicted pose lies outside the safety boundary, a new command is sampled. After a fixed number of retries, a fallback command that drives the robot toward the origin is used instead. To mitigate odometry drift, the safety region can be re-centered around the robot’s current position using a joystick override. In total, we collect approximately one hour of data corresponding to 180K state transitions.
Figure S6: RViz visualization of the real-world data collection process. The green arrow indicates a sampled base velocity command; the red sphere marks the estimated final pose via finite integration. The red box illustrates the safety boundary around the odometry origin, which the robot must remain within during data collection.
A.4.3 Additional Offline Data Sources
RWM-O can achieve reasonable performance when trained exclusively on real-world data, largely due to the inherent distributional bias in the dataset. Since all trajectories are collected from an expert policy without failures such as falls or collisions, the world model only needs to generalize within a narrow, well-behaved region of the state-action space. MOPO-PPO capitalizes on this bias by penalizing transitions with high epistemic uncertainty, discouraging the policy from entering regions outside the dataset’s support. However, this mechanism alone is insufficient for fully stable policy learning. In particular, the absence of privileged signals, such as contact labels, limits the ability of RWM-O to predict physically grounded dynamics, especially during transitions involving contact events. These signals are critical for accurate reward computation, and their absence introduces noise in policy gradients, leading to instability during training. To mitigate this, additional offline data from simulation, with access to privileged information, can be incorporated to enhance the world model’s predictive capabilities. This improves the fidelity of reward estimation during imagination rollouts, resulting in more stable and robust policies that transfer reliably to hardware.